OpenStudy (anonymous):

Solve the system using elimination: 2x + y = 20 6x - 5y = 12

OpenStudy (anonymous):

5(2x+y=20) 10x=5y=120 6x-5y=12 4x= 108 x=27

OpenStudy (anonymous):

6(27)-5y=12 162-5y=12 -5y=-150 y=30

OpenStudy (anonymous):

(27,30)

OpenStudy (anonymous):

you want to match up the constants for either the x, the y, or the singular constant. That way they'll cancel out For example, if you multiplied 2x+y=20 by 5, both equations would have 5y which would then cancel out when you add them since one 5y is negative. That would leave you with just x's which you could solve for. Once you get the x value, you can plug it into either equation and get the answer for y

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (anonymous):

No problem :)